OpenStudy (anonymous):

f(x)=7/8|X+3|-9 Graph and tell what the following are: Does it open up or down Wider,narroweer,or same vertex?

OpenStudy (anonymous):

Which one are you having trouble with?

OpenStudy (anonymous):

the vertex?

OpenStudy (anonymous):

Remember when you have f(x)=a|x-h|+k, the vertex is (h,k)

OpenStudy (anonymous):

i know but i am still having trouble

OpenStudy (anonymous):

do you know what k is?

OpenStudy (anonymous):

no nither H

OpenStudy (anonymous):

f(x)=a|x-h|+k f(x)=a|x+3|-9 What if i place them like this? Note that a=7/8 on the second one. Removed so they line up better

OpenStudy (anonymous):

okay so H is -9 right?

OpenStudy (anonymous):

and k is 7/8

OpenStudy (anonymous):

Look again and see which one goes with the h and which one goes with the k. They are lined up on top of each other

OpenStudy (anonymous):

oh h is 3 and k is -9

OpenStudy (anonymous):

Be careful with the h. Notice the sign change. You can solve it by equating whats inside the abs So, x-h=x+3.

OpenStudy (anonymous):

so its -3 and -9 and it's narrower and opens up?

OpenStudy (anonymous):

It opens up but its not narrower. See |a|<1 where a=(7/8). So it's wider

OpenStudy (anonymous):

oh okay. is my vertex right -3,-9

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

okay thank you.

OpenStudy (anonymous):

Np